OpenStudy (anonymous):

the sum of numberand its reciprocal is 73/24. find the number?

OpenStudy (anonymous):

let the number be x. so, \[x+{1\over{x}}={73\over24}\]

OpenStudy (anonymous):

\[{x^2+1\over{x}}={73\over24}\]

OpenStudy (anonymous):

\[24x^2-73x+24=0\] \[\left( x-{8\over3} \right)\left( x-{3\over8} \right)=0\] \[x={8\over3},{3\over8}\] so just choose either answer (: